Shift summary for day shift, 17/Dec/99

AG (
Fri, 17 Dec 1999 16:57:43 -0500 (EST)

Shift Summary for day shift, 17/Dec/99

On Shift: DJ, AG

Hours of beam available at all intensities: 7.2h
Hours of beam available at nominal intensity: 7.2h
Hours of E799 data (special runs not included): 7.2h
Hours of special runs (laser scans, muon runs...): 0.0h

Accounting of lost time:

o couple of spills here and there.
o Repairs to MI-60L started at 1420.


Runs ended this shift:
Run 15237 - 25.4M - BEAM_799

Tapes brought to FCC this shift:
(including tapes for on-line split)



Accesses during shift:

1. DJ and AG went on access, so that AG could change aerogel setup
configuration and drop the interlocks.
2. Search and secure to lock out the experimental hall.


Overview of shift:

Very smooth E799 data-taking until beam goes away at 1430 due to the
access to MI to work on a power supply.


Description of Shift:

run spill time comment


15237_________0700_______Day shift begins. Run 15237 in progress.

______________1320_______MCR phones to tell us that they plan an access to
_________________________the MI. It will start @ ~ 1420 and supposed to
_________________________last for ~ .5 hours.

______________1330_______Rick Ford phones to tell that he adjusted SWIC
_________________________HV, so the profiles do not saturate. Also
_________________________NM2 SEED's are in the beam now.

______________1430_______Beam's off.

15237_________1535_______Stop the run because KTeV DAQ monitoring window
_________________________died (actually, that happened at 12:11). Run
_________________________doesn't stop gracefully. Fresh start dart.

15238_________1556_______Started the run. MCR calls - 15 minutes till


Follow up Notes on Previous Shift


Items to Follow up on Next Shift